If cos(x+30)= sin x, a measure of angle x is ...
Answer by Alan3354(69443)

cos(x+30)= sin x
--------------
cos(x)*cos(30) - sin(x)sin(30) = sin(x)
cos(x)*sqrt(3)/2 - sin(x)/2 = sin(x)
cos(x)*sqrt(3)/2 = 3sin(x)/2
cos(x)*sqrt(3) = 3sin(x)
3cos^2(x) = 9sin^2(x)
cos^2(x) = 3sin^2(x)
3sin^2(x) = 1 - sin^2(x)
4sin^2(x) = 1
sin(x) = -1/2, 1/2
------------
You can finish and check it.